OpenStudy (anonymous):

Find the derivative of f(x)=[18e ^{0.5t}+2\]

OpenStudy (abb0t):

\[\frac{ d }{ dt }e^f(t) = e^{f(t)} \times \frac{ d }{ dt }f(t)\]

OpenStudy (abb0t):

chain rule

OpenStudy (anonymous):

\[f'(x)=9(e^.5)^t\]

OpenStudy (anonymous):

9^0=0